Circular No. 5959

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRS 1915+105 S. Sazonov, R. Sunyaev, and I. Lapshov, on behalf of the Granat

WATCH team, report: “The x-ray transient GRS 1915+105, discovered by the WATCH instrument on 1992 Aug. 15 (IAUC 5590), appeared bright in recent WATCH observations: the fluxes in the band 8–20 keV for Mar. 14, 18, 21, 24, and 27 were 380 ± 60, 400 ± 90, 220 ± 40, 330 ± 80, and 180 ± 30 mCrab, respectively. The previous WATCH detection of GRS 1915+105 in its high state occured at the end of May 1993.”

GRO J1008–57 M. H. Finger, R. B. Wilson, M. Scott, M. Stollberg, and T. A. Prince

report for for the Compton Gamma Ray Observatory (CGRO) BATSE pulsar team: “Pulsed hard x-ray flux in the energy range 20–100 keV is currently being detected from the source GRO J1008–57. This source was discovered by CGRO in an outburst in 1993 July (IAUC 5836, 5838) and subsequently observed by ASCA (IAUC 5851) and ROSAT (IAUC 5877). In the current outburst, pulsed flux was first detected on 1994 Mar. 23. The barycentric pulse period of 93.5410 ± 0.0014 s was determined for Mar. 27.0. The pulse profile consists of a single broad peak, as was previously observed by BATSE. The spectrum of the phase-averaged pulsed flux on Mar. 26 is fit between 20 and 100 keV by an optically-thin thermal bremsstrahlung model of the form A/E exp(−E/kT), with a flux at 45 keV of (1.0 ± 0.2) × 10^{−4} photon cm^{−2} s^{−1}, and kT = 17.9 ± 4.3 keV. This is ∼ 40 mCrab (total) in the band 20–50 keV. If the interval between this outburst and the last (260 days) is near the orbital period, then a time-of-arrival analysis of the previous outburst requires a projected semi-major axis of 600-800 light seconds, and a mass function of 3–8 M_{⊙}, indicating a highmass companion.”

SUPERNOVA 1994D IN NGC 4526 Visual magnitude estimates (cf. IAUC 5953): Mar. 18.023 UT, 11.9

(B. H. Granslo, Fjellhamar, Norway); 19.02, 11.8 (B. H. Granslo, Fjellhamar, Norway); 19.957, 12.0 (G. M. Hurst, Basingstoke, England); 22.01, 11.8 (Granslo); 25.657, 11.6 (S. Takahashi, Shiga, Japan); 27.032, 11.8 (Granslo); 28.014, 11.9 (H. Dahle, Oslo, Norway).

1994 March 29 Daniel W. E. Green